The number of people attending a concert is represented by 9000(75-x), where x is the cost of the ticket in dollars. What ticket price will maximize the total revenue of the concert?


1
Expert's answer
2021-03-26T08:14:50-0400

Given, Price of ticket is"\" "x" and number of people attending the concert are "\ 9000(75-x)" .

So ,total revenue generated will be "R=x\times 9000(75-x)=9000x(75-x)"

To maximize "\ R,\frac{dR}{dx}=0" and "\ \frac{d^2R}{dx^2}<0"

So, "\frac{dR}{dx}=\frac{d}{dx}(9000x(75-x))=9000[x(-1)+(1)(75-x)]=9000[75-2x]"

And "\frac{dR}{dx}=0\implies 9000[75-2x]=0\implies x=\frac{75}{2}=37.5"

Also ,"\frac{d^2R}{dx^2}=9000(-2)=-18000<0"

So, for "x= 37.5 ," R is maximum.

Hence,  ticket price will maximize the total revenue of the concert is "37.5" dollars.